OpenStudy (carlyleukhardt):

PLEASE HELP!! MEDAL AND FAN Add 1/10 + 11/100 and select the correct answer. A) 1 100 B) 1 5 C) 11 100 D) 21 100

OpenStudy (anonymous):

a

OpenStudy (carlyleukhardt):

thanks!!!

OpenStudy (anonymous):

no problom

OpenStudy (anonymous):

reported, no work was shown

OpenStudy (carlyleukhardt):

stop @jacobrocket173

OpenStudy (anonymous):

so

OpenStudy (anonymous):

what, hes not suppose to say the answer especially plus he didn show how to get his answer

OpenStudy (anonymous):

its against the code of conduct, im sry but i have to report you

OpenStudy (anonymous):

plus your actually wrong

OpenStudy (carlyleukhardt):

Who cares, hes awesome!

OpenStudy (carlyleukhardt):

Jacob be a good bf please stop

OpenStudy (anonymous):

1/10 turns to 10/100 10/100+11/100=21/100

OpenStudy (carlyleukhardt):

Oh

OpenStudy (carlyleukhardt):

thanks <3333333333

OpenStudy (anonymous):

that is why you dont trust someone that says the answer right off the bat like that and report them...

OpenStudy (anonymous):

sry

OpenStudy (carlyleukhardt):

Which expression is equivalent to (2 + 3) + 5? A) 2 + (3 + 5) B) 2(3 + 5) Eliminate C) 2 + 15 D) 6 + 5

OpenStudy (carlyleukhardt):

help me with this one babe

OpenStudy (anonymous):

well, the comulative prop. says that a+(b+c)=(a+b)+c

OpenStudy (carlyleukhardt):

what?

OpenStudy (carlyleukhardt):

just answer the question

OpenStudy (anonymous):

based from the comulative prop. its A

OpenStudy (carlyleukhardt):

babe your the best <3

OpenStudy (anonymous):

thx baby

OpenStudy (carlyleukhardt):

What is the sum of 14 and the opposite of 12?

OpenStudy (anonymous):

well, 14+(-12) 14-12=2

OpenStudy (carlyleukhardt):

You are so sweet <3 u deserve 20000 medals

OpenStudy (anonymous):

awwe :$

OpenStudy (carlyleukhardt):

I love u <3 can u help with like 2 more?

OpenStudy (anonymous):

sure, if i still have internet connection

OpenStudy (anonymous):

you dont like me :(

OpenStudy (carlyleukhardt):

What is the distance between -45 and -98 on a number line? A) -67 B) 53 C) 67 D) 133

OpenStudy (anonymous):

am i a screwup

OpenStudy (carlyleukhardt):

@LegendaryClasher awh yes i do your cool

OpenStudy (carlyleukhardt):

no your not!!! :)

OpenStudy (anonymous):

well you did have the wrong answer and you didnt show how you got that answer but that doesnt mean that your better at something else

OpenStudy (carlyleukhardt):

babe help with what i just asked>.<

OpenStudy (anonymous):

YOU REPORTED ME!!!!!!!!